If the sides of a square are lengthened by 8cm, the area becomes 144cm^2
find the length of a side of the original square.
I could use some help with this.

Found 2 solutions by nyc_function, scott8148:
Answer by nyc_function(2741)

Let x = side of original square.
Let x + 8 = side of new square.
The area is found using A = side times side.
(x + 8)^2 = 144
Can you finish now?
All you need to do is find x.

Answer by scott8148(6628)

(x + 8)^2 = 144

x + 8 = 12

x = 4